OpenStudy (anonymous):

(6+3√2) (5-2√2) solve?

myininaya (myininaya):

\[6(5-2 \sqrt{2})+3 \sqrt{2}(5- 2 \sqrt{2})=6(5)+6(-2 \sqrt{2})+3\sqrt{2}(5)+3 \sqrt{2}(-2 \sqrt{2})\]

myininaya (myininaya):

\[30+(-12 \sqrt{2})+15 \sqrt{2}+(-6\cdot2)=30-12 \sqrt{2}+15 \sqrt{2}-12=18+3 \sqrt{2}\]

OpenStudy (anonymous):

thanks